Polymer Industries uses weighted average process costing to value its inventories. For May, it furnishes the following information regarding the final step (the finishing department) in making its product. This department receives the product from the mixing department.

• The only additional cost incurred in the finishing department is conversion, which is incurred uniformly.

• Opening WIP of 45,000 units was valued at $505,000 for transferred in costs and $41,600 for conversion. These units were 20% done for conversion.

• The finishing department received 155,000 units during May. The department’s WIP account was debited $1,745,000 toward the transferred in costs of these units.

• The department spent $762,000 toward conversion costs.

• 140,000 units were completed and transferred out during May.

• Ending WIP was 40% done for conversion.

Determine the following (compute cost / EU to four decimal places, if needed):

Number of units in ending WIP    Value of ending WIP
Cost per equivalent unit for May (transfer in) Value of units transferred out (COGM)   
Cost per equivalent unit for May (conversion)
Answer #1
1.) Number of units in Ending WIP              60,000 =45000+155000-140000
2.) Transfer in
Units Completed & Transfer out           140,000
Ending WIP -60,000              60,000
Equivalent Units          200,000
Transfer in
Cost of Beginning WIP           505,000
Add: Cost of transferred units        1,745,000
Total Cost $        2,250,000
Cost per Equivalent unit for May $ 11.2500 =2250000/200000
( Transfer in )
3.) Conversion
Units Completed & Transfer out           140,000
Ending WIP -60,000 x 40%              24,000
Equivalent Units          164,000
Conversion
Cost of Beginning WIP              41,600
Add: Cost of conversion of May           762,000
Total Cost $          803,600
Cost per Equivalent unit for May $ 4.9000 =803600/164000
( Conversion )
4.) Amount $
Units Transfer In           675,000 =60000*11.2500
Conversion           117,600 =24000*4.9000
Value of Ending WIP          792,600
5.) Amount $
Units Transfer In        1,575,000 =140000*11.2500
Conversion           686,000 =140000*4.9000
Value of units transferred out        2,261,000